# Multi-Grained Query-Guided Set Prediction Network for Grounded Multimodal Named Entity Recognition

### v1

@@ 3.6. Multimodal Set Prediction Network
Twitter-GMNER Twitter-FMNERG Train Dev Test Train Dev Test #Tweet 7000 1500 1500 7000 1500 1500 #Entity 11,782 2,453 2,543 11,779 2,450 2,543 #Groundable Entity 4,694 986 1,036 4,733 991 1,046 #Box 5,680 1,166 1,244 5,723 1,171 1,254
4. Experiments
4.1. Experiment Settings
@@
GMNER Datasets. We conduct experiments on the Twitter-GMNER (Yu et al., 2023) and Twitter-FMNERG (Wang et al., 2023). Twitter-GMNER only contains four entity types: Person (PER), Organization (ORG), Location (LOC), and Others (OTHER) for text-image pairs. Twitter-FMNERG extends the GMNER dataset to 8 coarse-grained and 51 fine-grained entity types. Both of them are built based on two publicly MNER Twitter datasets, i.e., Twitter-2015 (Zhang et al., 2018) and Twitter-2017 (Lu et al., 2018a). Table 1 shows the details of Twitter-GMNER and Twitter-FMNERG.
Evaluation Metrics. The GMNER task can be decomposed into multimodal entity recognition(MNER) and entity extraction and grounding(EEG). Following (Yu et al., 2023), The correctness of each prediction is computed as follows:
(17) $C_{e}/C_{t}=\begin{cases}1,&p_{e}/p_{t}=g_{e}/g_{t};\\ 0,&\text{otherwise.}\end{cases}$
(18) $C_{r}=\begin{cases}1,&p_{r}=g_{r}=None;\\ 1,&\max(\mathrm{IoU}_{1},...,\mathrm{IoU}_{j})>0.5;\\ 0,&\text{otherwise.}\end{cases}$
@@
(19) $correct$ =$=\begin{cases}1,\quad C_{e}\mathrm{~{}and~{}}C_{t}\mathrm{~{}and~{}}C_{o};\\ 0,\quad\mathrm{otherwise}.\end{cases}$
$Pre=\frac{\#correct}{\#predict},\quad Rec$ =#$=\frac{\#correct}{\#gold},\quad F1=\frac{2\times Pre\times Rec}{Pre+Rec},$
@@
Implementation Details All experiments are implemented on NVIDIA RTX3090 GPU with Pytorch 1.8.0. For a fair comparison, we use the pre-trained BERT-based model (Devlin et al., 2019) as the textual encoder, ViT-B/32 (Dosovitskiy et al., 2020) as the visual encoder, and VinVL (Zhang et al., 2021a) as a class-agnostic RPN. We assign the layer of the QFNet as $L=3$ and initialize the learnable query part with a normal distribution $\mathcal{N}(0,0.02)$. For training, we set the batch size to 16, the learning rate to $2\times 10^{-5}$, and the training epoch to 50. Our model uses an Adam optimizer with a linear warmup of ratio 0.05. We tune the optimal query number $u$ to 60 and 120 for Twitter-GMNER and Twitter-FMNERG datasets, respectively. To allow the multi-grained queries to learn query semantics initially, we first train the model for 5 epochs with the freezing parameters of BERT and ViT.
4.2. Baselines
@@
We categorize the existing baseline methods into three groups: (i) Text-only methods, (ii) Text+Image pipeline methods, and (iii) Unified methods.
(i). The first group first uses representative text-based NER models to extract the corresponding entities and then sets the region prediction to the majority class, i.e., None. We consider the following text-based NER models: (1) Sequence-labeling methods: HBiLSTM-CRF (Lu et al., 2018b) and BERT-CRF (Devlin et al., 2019). (2) MRC-based method: BERT-MRC (Li et al., 2020). (3) Sequence generation-based method: BARTNER (Yan et al., 2021). (4) Set prediction-based method: Seq2Set (Tan et al., 2021).
@@
(ii). The second group first uses any previous state-of-the-art MNER method and object detector, i.e., VinVL (Zhang et al., 2021a) or Faster R-CNN (Girshick, 2015), to pre-extract entity-type pairs and candidate regions. Entity-aware Visual Grounding (EVG) module (Yu et al., 2023) is then used to predict the matching relationship between entities and regions. The following MNER models are utilized as strong baselines: (1) GVATT (Lu et al., 2018b) adopts a visual attention mechanism based on BiLSTM-CRF to extract multimodal entities. (2) UMT (Yu et al., 2020) proposes a multimodal transformer to capture the cross-modality semantics. (3) UMGF (Zhang et al., 2021c) solves text-image integration with a multimodal graph fusion mechanism. (4) ITA (Wang et al., 2022b) utilizes the image-text translation and object tags to explicitly align visual and textual features. (5) BARTMNER (Yu et al., 2023) expands generative model BART with cross-modal transformer layer.
@@
(iii). The third group includes several unified GMNER approaches. (1) H-Index (Yu et al., 2023) uses a multimodal BART model with the pointer mechanism to formulate the GMNER task as a sequence generation. (2) TIGER (Wang et al., 2023) is T5-based generative model which converts all span-type-region triples into target paraphrase sequences. (3) MNER-QG (Jia et al., 2023) is a unified MRC framework that combines entity extraction and grounding with multi-task joint learning.
@@
For a fair comparison, we did not compare with methods using large language models (Li et al., 2024) or external knowledge-enhanced (Ok et al., 2024).
Table 2. Performance comparison of different competitive approaches on Twitter-GMNER datasets. Bold represents the optimal result, and underlined represents the suboptimal result. For the baseline methods, $\clubsuit$ indicates that the results are obtained by running the code released by the corresponding papers, and others are from (Yu et al., 2023).
[…]
Table 3. F1 scores of different methods on Twitter-FMNERG datasets for fine-grained entity type. Baseline results are from (Wang et al., 2023).
Methods Person Location Building Organization Product Art Event Other All GVATT-VinVL-EVG 35.21 61.64 35.37 42.60 15.38 32.79 42.55 41.03 40.32 UMT-VinVL-EVG 37.10 63.58 35.09 42.82 18.28 36.62 46.79 38.24 41.32 UMGF-VinVL-EVG 37.04 63.16 38.51 44.71 17.39 30.70 47.81 38.89 41.92 ITA-VinVL-EVG 37.91 65.52 39.16 44.34 17.18 34.91 46.62 36.36 42.78 H-Index 45.13 62.33 32.88 46.68 28.19 38.89 45.56 41.81 46.55 TIGER 43.78 67.69 40.00 46.75 27.38 43.27 48.39 48.28 47.20 MQSPN(Ours) 54.75 75.63 47.56 56.19 29.02 42.99 55.02 39.10 50.03
4.3. Overall Performance
Performance on GMNER, MNER, and EEG. Following (Yu et al., 2023), we also report two subtasks of GMNER, i.e., Multimodal Named Entity Recognition(MNER) and Entity Extraction & Grounding(EEG). MNER aims to identify entity-type pairs, while EEG aims to extract entity-region pairs. Table 2 shows the performance comparison of our method with baseline models on twitter-GMNER benchmarks.
First, Seq2Set-None can obtain comparable F1 scores in text-only methods, which underscores the potential of set prediction-based approaches in GMNER tasks. Second, unified models are significantly superior to pipeline methods due to joint training of MNER and EEG to mitigate error propagation. Third, compared with the previous state-of-the-art unified model H-index (Yu et al., 2023), our method MQSPN exhibits superior performance, achieving $+0.7\%$, $+1.22\%$, and $+1.11\%$ F1 scores improvements on the MNER, EEG, and GMNER tasks, respectively. Furthermore, MQSPN significantly outperforms MNER-QG by $+3.61\%$ F1 scores.
We attribute the performance improvements of MQSPN to the following reasons: (1) Compared with the sequence generation-based approach H-Index and TIGER, MQSPN eliminates the dependencies on predefined decoding order and establishes suitable inter-entity relationships from a global matching view. (2) Compared with the MRC-based approach MNER-QG, MQSPN can learn fine-grained query semantics and model intra-entity connections between regions and entities.
Performance on fine-grained GMNER. We validate the fine-grained GMNER ability of different methods on Twitter-FMNERG datasets, and the experimental results are shown in Table 3. It reveals that our proposed MQSPN achieves the best results on all fine-grained entity types except Other and Art among the state-of-the-art models. Specifically, we achieve a great improvement of $+9.62\%$, $+7.94\%$, $+7.56\%$, $+9.44\%$, $+6.63\%$, and $+2.83\%$ F1 scores on the Person, Location, Building, Organization, Event, and All entity types respectively. The experimental results demonstrate that compared to the SOTA sequence generation-basd method TIGER, our MQSPN exhibits superior capability in modeling fine-grained GMNER task.
4.4. Ablation Study
Ablation Setting. To verify the effectiveness of each designed component in MQSPN, we conduct a comprehensive ablation study on the Twitter-GMNER and Twitter-FMNERG dataset: For Multi-grained Query Set (MQS), (1) w/o POQ: we replace the prompt-based type-grained part of the query set with randomly initialized embedding. (2) w/o LEQ: we remove the learnable entity-grained part in query construction. For Query-guided Fusion Net (QFNet), (3) w/o QCT: we encode sentences and the query set using the original BERT without query-text cross-attention. (4) w/o QPI: we eliminate the query-region prefix integration. (5) w/o SAG: we eliminate the similarity-aware aggregator. For Multimodal Set Prediction (MSP), (6) w/o BML: we replace the bipartite matching loss with a joint cross-entropy loss in the fixed permutation of the entities. The experimental results are shown in Table 4.
Multi-grained Query Set. A notable contribution of MQSPN is the design of a Multi-grained Query Set (MQS), which includes a prompt-based type-grained part and a learnable entity-grained part. In Table 4, we observe a clear F1 scores drop in model performance (1.57% in Twitter-GMNER and 1.08% in Twitter-FMNERG) without the prompt-based type-grained queries, indicating that integrating type-grained information into the query can enhance model performance. Furthermore, adding an additional learnable entity-grained part to the query substantially improves model performance: in the Twitter-GMNER and Twitter-FMNERG datasets, the F1 scores increased by $+2.50\%$ and $+2.82\%$, respectively. The main reason is that these learnable vectors can effectively learn the entity-grained semantics and model intra-entity connections between entity spans and regions, thereby boosting model performance.
Figure 3. Analysis of the multi-grained queries quantity $u$ on GMNER, MNER and EEG tasks.
Table 4. Ablation study of each component on overall F1 score of different tasks.
Module Settings Twitter-GMNER Twitter-FMNERG GMNER MNER EEG GMNER MNER EEG MQSPN Default 57.52 80.43 62.40 50.03 70.16 62.58 MQS w/o PTQ 55.95 ($\downarrow$1.57) 78.78 60.51 48.95 ($\downarrow$1.08) 68.78 61.94 w/o LEQ 55.02 ($\downarrow$2.50) 77.64 59.30 47.21($\downarrow$2.82) 67.06 59.27 QFNet w/o QCT 56.21 ($\downarrow$1.31) 79.34 60.88 48.94($\downarrow$1.09) 68.71 61.93 w/o QPI 56.86 ($\downarrow$0.66) 79.81 61.51 49.64($\downarrow$0.39) 69.69 62.29 w/o SAG 57.13 ($\downarrow$0.39) 79.73 62.06 49.16($\downarrow$0.87) 69.20 61.97 MSP w/o BML 55.58 ($\downarrow$1.94) 78.26 60.37 48.39($\downarrow$1.64) 68.34 60.21
Query-guided Fusion Net. We investigate the impact of the Query-guided Fusion Net (QFNet), which comprises query-text cross-attention (QCT), query-region prefix integration (QPI), and similarity-aware aggregator (SAG). The model without these three modules achieves consistent drops. We find the average F1 scores of these three modules decreased by $0.79\%$ and $0.78\%$ on Twitter-GMNER and Twitter-FMNERG, respectively. Experimental results demonstrate that the QFNet can fuse information across different modalities and filter the noise introduced by irrelevant visual regions, which is crucial for improving the performance of MQSPN.
Bipartite Matching Loss. We evaluate the effect of bipartite matching, which is used in Multimodal Set Prediction (MSP) module to ensure the optimal matching between predictions and ground truths. Excluding the bipartite matching loss from the model leads to an F1 scores decline of $1.94\%$ and $1.64\%$ in Twitter-GMNER and Twitter-FMNERG datasets. This illustrates that the bipartite matching loss contributes to model performance through modeling inter-entity relationships in global matching.
Figure 4. Analysis of candidate visual regions number $k$ for H-Index (Yu et al., 2023), TIGER (Wang et al., 2023), and our MQSPN.
Figure 5. Predictions of H-Index and our MQSPN on two test samples. The red part in prediction means the occurrence of error
4.5. Discussion and Analysis
Analysis of multi-grained queries quantity. To explore the impact of multi-grained query quantity (i.e., $u$) on GMNER and its 2 subtasks, we report the F1 scores of Twitter-GMNER in Figure 3 by tuning $u$ from $20$ to $90$. We observe that as the number of queries increases from $20$ to $60$, the model’s performance improves by $+1.66\%$, $+1.76\%$, and $+1.45\%$ in the GMNER, MNER, and EEG tasks, respectively. These results indicate that query quantity plays a crucial role in intra-entity connection learning. However, a query quantity exceeding $60$ does not lead to better performance. This suggests that there is an optimal number of queries for MQSPN. In our experiments, we set $u=60$ to achieve the optimal results.
@@ 4.5. Discussion and Analysis
Table 5. Analysis of impact on QFNet layers on Twitter-GMNER.
#Layers GMNER Pre. Rec. F1. 1 57.32 55.70 56.50 2 57.61 56.01 56.80 3 58.21 56.84 57.52 4 58.15 56.77 57.45 5 58.30 56.87 57.57
@@
Analysis of Query-guided Fusion Net Layers. We analyze the influence of the number of QFNet layers on the performance of MQSPN. As presented in Table 5, when the number of fusion layers is decreased from 3 to 2 and from 2 to 1, the overall F1 scores drop by $0.72\%$ and $0.30\%$, respectively. Furthermore, increasing the number of fusion layers from 3 to 5 results in only a marginal performance improvement of $0.05\%$; however, it comes at the cost of decreased inference speed in the model. It also reveals that MQSPN is insensitive to the layers of the cross-modal fusion Layers.
Sensitivity Analysis of Irrelevant Visual Regions. The performance of EEG is largely determined by the ground truth coverage of the top-$k$ regions proposed by the class-agnostic RPN. However, a higher value of $k$ implies the introduction of more irrelevant visual regions. To delve into the model’s error tolerance to noisy visual regions, we conduct experiments on H-Index, TIGER, and MQSPN under different values of $k$. As shown in Figure 4, we can observe that the F1 scores of these three methods first increase, and then decrease with the augment of $k$. Compared with H-Index and TIGER, our proposed MQSPN has better and more stable performance on both EEG and GMNER tasks when $k$ becomes larger. This demonstrates MQSPN’s insensitivity to irrelevant visual regions, indicating its excellent error tolerance and robustness.
Case Study. To better understand the advantage of our proposed approach, we compare the predictions of the previous SOTA method H-Index (Yu et al., 2023) and MQSPN on two test samples in Twitter-GMNER dataset. In Figure 5 (a), H-Index incorrectly identifies the entangled entity ”Jimmy Kimmel” within ”Jimmy Kimmel Live” as PER, and associates it with the blue box. This results in incorrectly matching the azure box with ”The Weeknd”, which originally belongs to the blue box. In contrast, our MQSPN accurately outputs these two PER entities with their corresponding bounding boxes and identifies ”Jimmy Kimmel Live” as OTHER without a box. In Figure 5 (b), both two methods correctly identify 3 textual PER entities. However, H-Index misunderstands the relation between three PER entities and two boxes, which wrongly aligns ”Donald Trump Jr” with a blue box and ”Stephen Colbert” with the azure box. Conversely, MQSPN can establish correct connections between all PER entities and boxes. It illustrates that MQSPN can learn distinguishable semantics to model relationships between same-type entities and regions.
Conclusion
In this paper, we propose a novel unified framework named MQSPN that consists of Multi-grained Queries Set (MQS), Multimodal Set Prediction Network (MSP), and Query-guided Fusion Net (QFNet). The three modules cooperate to model appropriate intra-entity and inter-entity relationships for the GMNER task. Compared with MRC-based methods, MQS provides learnable entity-grained queries to adaptively model solid connections between entity spans and regions, mitigating intra-entity relationship misunderstanding. Compared with sequence generation-based methods, MSP reformulates GMNER as a set prediction to eliminate inter-entity overreliance caused by predetermined sequence decoding order. Experimental results demonstrate that our MQSPN achieves state-of-the-art performance on both GMNER and its 2 subtasks across two Twitter benchmarks. Comprehensive ablation studies verify the effectiveness of each designed module. In the future, the extension of our MQSPN framework to the vision-language pre-training model will be one of our directions.

### v2

@@ Conclusion
Twitter-GMNER Twitter-FMNERG Train Dev Test Train Dev Test #Entity type 4 4 4 51 51 51 #Tweet 7000 1500 1500 7000 1500 1500 #Entity 11,782 2,453 2,543 11,779 2,450 2,543 #Groundable Entity 4,694 986 1,036 4,733 991 1,046 #Box 5,680 1,166 1,244 5,723 1,171 1,254
Table 5: The statistics of two GMNER Twitter datasets.
@@
In this work, we conduct experiments in 2 tweet datasets, Twitter-GMNER and Twitter-FMNERG. Twitter-GMNER only contains four entity types: Person (PER), Organization (ORG), Location (LOC), and Others (OTHER) for text-image pairs. Twitter-FMNERG extends the GMNER dataset to 8 coarse-grained and 51 fine-grained entity types. Both of them are built based on two publicly MNER Twitter datasets, i.e., Twitter-2015 (Zhang et al. 2018) and Twitter-2017 (Lu et al. 2018a). Table 5 shows the statistical details of Twitter-GMNER and Twitter-FMNERG.
Evaluation Metrics.
The GMNER prediction is composed of entity span, type, and visual region. Following previous works(Yu et al. 2023), The correctness of each prediction is computed as follows:
$C_{e}/C_{t}=\begin{cases}1,&p_{e}/p_{t}=g_{e}/g_{t};\\ 0,&\text{otherwise.}\end{cases}$ (16)
$C_{r}=\begin{cases}1,&p_{r}=g_{r}=None;\\ 1,&\max(\mathrm{IoU}_{1},...,\mathrm{IoU}_{j})>0.5;\\ 0,&\text{otherwise.}\end{cases}$ (17)
@@
$correct=\begin{cases}1,\quad C_{e}\mathrm{~{}and~{}}C_{t}\mathrm{~{}and~{}}C_{o};\\ 0,\quad\mathrm{otherwise}.\end{cases}$ (18)
$Pre=\frac{\#correct}{\#predict},\quad Rec$ =#$=\frac{\#correct}{\#gold},$
$\quad F1=\frac{2\times Pre\times Rec}{Pre+Rec}$
@@
Baselines
@@
We categorize the existing baseline methods into two groups: text+image pipeline methods and unified methods.
@@
The first group first uses any previous state-of-the-art MNER method and object detector, i.e., VinVL (Zhang et al. 2021c) or Faster R-CNN (Girshick 2015), to pre-extract entity span-type pairs and type-related candidate regions. Entity-aware Visual Grounding (EVG) module (Yu et al. 2023) is then used to predict the matching relationship between entities and regions. The following MNER models are utilized as strong baselines: (1) GVATT (Lu et al. 2018b) adopts a visual attention mechanism based on BiLSTM-CRF to extract multimodal entities. (2) UMT (Yu et al. 2020) proposes a multimodal transformer to capture the cross-modality semantics. (3) UMGF (Zhang et al. 2021a) solves text-image integration with a multimodal graph fusion mechanism. (4) ITA (Wang et al. 2022b) utilizes the image-text translation and object tags to explicitly align visual and textual features. (5) BARTMNER (Yu et al. 2023) expands generative model BART with cross-modal transformer layer.
@@
The second group includes several unified GMNER approaches. (1) H-Index (Yu et al. 2023) uses a multimodal BART model with the pointer mechanism to formulate the GMNER task as a sequence generation. (2) TIGER (Wang et al. 2023) is T5-based generative model which converts all span-type-region triples into target paraphrase sequences. (3) MNER-QG (Jia et al. 2023) is a unified MRC framework that combines textual entity extraction and grounding with multi-task joint learning. Please note that we replace the weakly-supervised query grounding in MNER-QG with a fully-supervised mode.
@@
For a fair comparison, we did not compare with methods using large language models (Li et al. 2024a) or knowledge-enhanced methods (Ok et al. 2024; Li et al. 2024b).
Analysis of Query-guided Fusion Net Layers
#QFNet Layers Twitter-GMNER Twitter-FMNERG Pre. Rec. F1. Pre. Rec. F1. 1 58.32 56.70 57.50 48.16 46.57 47.35 2 58.61 57.01 57.80 49.04 47.29 48.15 3 59.03 58.49 58.76 49.21 47.94 48.57 4 59.13 57.67 58.39 48.73 47.65 48.18 5 59.16 58.58 58.87 49.09 47.72 48.40
@@
Table 6: Analysis of the impact of QFNet layers in GMNER task across two datasets.
Figure 6: Comparison of inference speed on Twitter-GMNER and Twitter-FMNERG. All experiments are conducted on four NVIDIA RTX3090 GPUs with 24GB graphical memory.
Figure 7: Visualization of predictions in Twitter-FMNERG benchmark. These examples demonstrate our MQSPN can effectively establish intra-entity connections between entity spans and entity regions, and distinguish ambiguous entities such as Adidas (company) and Adidas Ace (brand product). Meanwhile, the appropriate modeling of inter-entity relationships by MQSPN renders current predictions insensitive to previous errors.
@@
We analyze the influence of the number of QFNet layers on the performance of MQSPN. As presented in Table 6, when the number of fusion layers is decreased from 3 to 2 and from 2 to 1, the overall F1 scores drop by $0.86\%$ and $0.30\%$ in Twitter-GMNER, and by $0.42\%$ and $0.80\%$ in Twitter-FMNERG, respectively. Furthermore, increasing the number of fusion layers from 3 to 5 results in only a marginal performance change; however, it comes at the cost of decreased inference speed in the model. It also reveals that MQSPN is insensitive to the layers of the QFNet.
Analysis of Inference Speed
We compare the inference speed on Twitter-GMNER and Twitter-FMNERG, as shown in Figure 6. The unified modeling approach for GMNER significantly accelerates inference speed compared to the pipeline method BARTMNER-VinVL-EVG. Notably, our MQSPN not only achieves superior performance but also demonstrates faster inference speeds than both MNER-QG (Jia et al. 2023) and TIGER (Wang et al. 2023). MNER-QG requires a type-specific query for each entity, necessitating multiple inference processes. Besides, the sequence generation-based method TIGER follows the autoregressive paradigm to predict span-type-region triplets one by one, leading to slower inference speed. However, our approach parallelly outputs all entities with a one-time input of MQS, making the inference faster.
More Case Analysis of Fine-grained GMNER
We further conduct more case analysis to compare the predictions of TIGER (Wang et al. 2023) and MQSPN for challenging samples in the Twitter-FMNERG dataset. In Figure 7 (a), both MQSPN and TIGER accurately recognize Calle Jarnkrok as an athlete and assign the correct entity region. However, TIGER incorrectly classifies Pred as other and fails to detect its corresponding region. In contrast, MQSPN correctly predicts both entities, even in the challenging case involving the sports team logo Pred. In Figure 7 (b), while TIGER recognizes Adidas, it fails to predict any span-type-region triplets for Adidas Ace, despite its multiple appearances in the image. Conversely, MQSPN not only detects the entity textual information but also accurately localizes all corresponding regions for Adidas Ace. This demonstrates MQSPN’s superiority in learning intra-entity connections, even when each entity is associated with multiple regions.
Figure 7 (b) and (c) show more difficult cases. TIGER entirely outputs incorrect entity predictions, encompassing errors such as entity misclassification (E.T.), incorrect entity grounding (E.R.), erroneous span boundary predictions (E.S.), and either missing entities (M.T.) or producing extraneous entities (A.T.). Compared to TIGER, MQSPN is insensitive to previous errors. TIGER incorrectly classifies Google as a company, which subsequently leads to the misclassification of Gabriela Mistral as a businessman due to one-by-one decoding order in an autoregressive manner. However, even when MQSPN misidentifies Google, it still correctly recognizes Gabriela Mistral, thanks to its global matching approach that models appropriate inter-entity relationships.
Besides, MQSPN successfully predicts several challenging entities. For example, in case (b), it correctly classifies Rosemary Barasa as a businessman and accurately matches it to the corresponding region. In case (c), MQSPN successfully differentiates Google as a website and correctly identifies the author Gabriela Mistral. However, MQSPN also generates some erroneous predictions. For instance, it fails to detect the logo of JKF Ag within the image in case (c) and cannot identify the tag for Google in case (d). These results indicate that GMNER remains a highly challenging task, highlighting the potential for further exploration in this area.